s d 33% 7:57 PM A webwork.phys.yorku.ca/webwork2/PHYS1010/w 2 of the questions remain unanswered. (1 point) THE PROBLEM A charge of q 2.9 pC is at the center of a cube. The dimensions of the cube are L x L × L where L-7.6 cm. Calculate the electric flux through one face of the cube. PAPER SOLUTION Solve the problem on paper first, including all four IDEA steps You will become a better physicist that way! Have you finished your paper solution already? oyes INTERPRET DEVELOP Derive an algebraic expression for the electric flux through one face of the cube. In your answer, type E0 to represent p. EVALUATE ASSESS Note: In order to get credit for this problem all answers must he eet ㄥ
Answer #1

Use Gauss' law for this problem. The electric flux through the entire cube is:

\Phi_t = \frac{q}{\epsilon_o}

now since the charge is at the center and there are 6 faces for a cube, the electric flux through one face of the cube is:

\Phi_1 = \frac{\Phi_t}{6} = \frac{q}{6\epsilon_o}.

Note that this answer is independent of the dimensions of the box.